### v1

# Scenario-Based Curriculum Generation for Multi-Agent Autonomous Driving

Abstract
The automated generation of diverse and complex training scenarios has been an important ingredient in many complex learning tasks. Especially in real-world application domains, such as autonomous driving, auto-curriculum generation is considered vital for obtaining robust and general policies. However, crafting traffic scenarios with multiple, heterogeneous agents is typically considered as a tedious and time-consuming task, especially in more complex simulation environments. In our work, we introduce MATS-Gym, a Multi-Agent Traffic Scenario framework to train agents in CARLA, a high-fidelity driving simulator. MATS-Gym is a multi-agent training framework for autonomous driving that uses partial scenario specifications to generate traffic scenarios with variable numbers of agents. This paper unifies various existing approaches to traffic scenario description into a single training framework and demonstrates how it can be integrated with techniques from unsupervised environment design to automate the generation of adaptive auto-curricula. The code is available at https://github.com/AutonomousDrivingExaminer/mats-gym.
I Introduction
Figure 1: This multi-agent scenario illustrates an intersection where five vehicles navigate according to assigned routes, with three pedestrians observed on the sidewalk adjacent to the ego vehicle. A visual representation of the simulation is depicted in the center, while the scenario description from which the simulation parameters are sampled is described on the right. On the left, we depict other scenarios sampled from the same Scenic description.
Autonomous driving (AD) stands as a groundbreaking frontier in transportation technology, offering the promise of safer and more efficient roadways. However, achieving fully autonomous vehicles has been a long-standing goal for researchers across generations. As recent advancements enable the deployment of more vehicles in cities worldwide, extensive testing in a diverse range of scenarios becomes essential. For every kilometer driven by AD systems on real-world roads, they traverse several orders of magnitude more in simulation. Thus, access to high-fidelity simulation environments plays a crucial role in advancing the field of AD. Open-source simulation software like CARLA [1] has empowered researchers and practitioners to evaluate and train AD algorithms in realistic environments. Despite this, creating diverse and lifelike traffic scenarios, incorporating varying numbers of agents, remains a laborious and time-consuming task. To tackle this challenge, scenario specification languages like Scenic [2] and OpenSCENARIO [^1] https://www.asam.net/standards/detail/openscenario/ have been introduced. These scenario specification approaches streamline the process of defining logical traffic scenarios and are well-integrated into simulation environments such as CARLA. However, we observe that (1) infrastructure aiding the training and evaluation of AD stacks is scattered across numerous ad hoc implementations, (2) integration of scenario specifications and simulation environments lacks support for training in multi-agent systems and (3) limited support for approaches targeting open-ended curriculum generation. To address these shortcomings, we propose MATS-Gym, an open-source multi-agent training framework that leverages existing traffic scenario-specification approaches for CARLA. Additionally, we demonstrate how MATS-Gym can be used to implement novel auto-curriculum approaches [3] for generating open-ended traffic scenarios with adaptive difficulty. Hence, the key contributions of MATS-Gym are:
  1. 1.
    A multi-agent, scenario-based AD training and evaluation framework for the CARLA simulator.
  2. 2.
    Experimental evaluation to demonstrate MATS-Gym’s practicality for multi-agent learning and ability to generate auto-curricula from scenario descriptions.
  3. 3.
    Extensive comparison and classification in the context of AD simulation frameworks.
I-A Illustrative Scenario
In this scenario, as depicted in Figure 1, five vehicles converge on a four-way intersection, each following a predefined route toward individual goal locations. Additionally, three pedestrians traverse the sidewalk near the intersection. All vehicles are equipped with sensors, including cameras, odometry, and velocity sensors. Despite the simplicity of this setting, the potential variations are vast, depending on initial conditions such as the number and types of actors and their behaviors. Effective training needs exposure to diverse scenario variations for robust generalization. Moreover, meticulous selection of progressively complex challenges is crucial but requires expert knowledge. In the following, we introduce MATS-gym and showcase its effectiveness in generating scenarios of varying difficulty.
II Related Work
TABLE I: Comparison of various traffic scenario frameworks
Framework Scenario Specification Realism Traffic Types Training
Sampling Scriptable Sensors Visual Physics Highway Urban RL MARL
HighwayEnv [4] $\sim$ $\sim$
BARK [5] $\sim$ $\sim$
CommonRoads [6, 7] $\sim$
SMARTS [8] $\sim$ $\sim$
MetaDrive [9] $\sim$ $\sim$
DI Drive [10]
MACAD-Gym [11] $\sim$
MATS-Gym (ours)
II-A Training Frameworks for AD
We compare our approach to other frameworks commonly utilized for training and evaluating AD agents. We identified several popular frameworks and assess them based on several key aspects crucial for our use-case. These aspects include ease and expressivity of scenario specifications, simulation realism, supported traffic types, and integrated training interfaces. Table I summarizes the comparison.
HighwayEnv [4], BARK [5], SMARTS [8], and the CommonRoads Suite [6, 7] offer alternatives to full-fledged game-engine simulations like CARLA. They focus on modeling vehicle dynamics and motion planning, with CommonRoads providing more sophisticated multi-body dynamics compared to HighwayEnv and BARK. While simpler environments are computationally efficient, they lack complex sensor models and realistic observations. These frameworks mainly involve vehicles and lack pedestrian or object modeling found in urban traffic. BARK and CommonRoads allow scriptable scenarios, while HighwayEnv supports custom scenarios through code re-implementation. SMARTS allows to model urban traffic scenarios with traffic lights and traffic signs. However, since the focus is not on graphical realism, those scenarios can not reach the visual diversity of traffic scenes in CARLA, which is necessary to train and evaluate vision based AD stacks.
MetaDrive [9] is a training framework that builds on the Panda3D engine. It offers a wide range of features such as procedural road generation and incorporating scenarios from traffic datasets. However, scenarios can not be specified by behavioral building blocks or predefined sub-scenarios. Moreover, its engine limits the structural variety of environments that can be achieved in CARLA.
DI Drive Gym [10] and MACAD Gym [11] are both training frameworks designed around CARLA. DI Drive Gym focuses on single-agent autonomous driving systems and interfaces with ScenarioRunner scenarios, lacking support for training multi-agent systems or scenario generation via sampling. MACAD Gym, supports multi-agent training but offers limited scenario generation capabilities, allowing only manual specification of initial conditions.
In this comparison, we focus on relevant aspects for learning-based approaches in realistic urban traffic scenarios. We acknowledge our criteria may not be complete and there might be aspects that could be relevant in other use cases. However, we tried to cover a broad spectrum of frameworks that could be used for training and evaluation.
II-B Curriculum Learning
Generating safety-critical scenarios is an active area of research and has been addressed in numerous works in recent years [12]. However, training agents in complex environments requires presenting levels at the right difficulty, matching their capabilities to generate meaningful learning signals. Curriculum Learning (CL) and Auto CL (ACL) methods have emerged to progressively expose agents to more complex environments. While traditional CL involves manually designing curricula, ACL automates this process by dynamically presenting suitable scenarios based on agent performance. For an overview, we refer to [13].
Previous works also applied ACL to AD scenarios [14, 15]. A novel research direction in ACL is Unsupervised Environment Design (UED), where task creation itself is treated as an optimization problem. This is formalized through Underspecified Partially Observable Markov Decision Processes (UPOMDP), allowing for flexible reparameterization of environments during runtime. A UPOMDP is a tuple
$\mathcal{M}=\langle A,O,\Theta,{S}^{\mathcal{M}},{\mathcal{T}}^{\mathcal{M}},{\mathcal{I}}^{\mathcal{M}},{\mathcal{R}}^{\mathcal{M}},\gamma\rangle,$
which differs from a standard POMDP only in having a set of free parameters $\Theta$, which can be chosen at any time during runtime. Recent works have explored different algorithms for UED. Prioritized Level Replay (PLR) utilizes domain randomization to generate parameters and a level replay buffer to resample scenarios proportional to their regret, a measure of suboptimality [16]. PAIRED [17] optimizes the level generator via reinforcement learning (RL) to maximize regret, while REPAIRED [3] combines PLR with the level generation approach from PAIRED in a Dual Curriculum Design (DCD) framework. Other approaches use evolutionary algorithms to optimize the generator and include a level editor for environment modifications [18, 19].
III MATS-Gym: Architecture and Features
MATS-Gym is a multi-agent training and evaluation framework that allows to generate diverse traffic scenarios in CARLA, a high-fidelity traffic simulator. The framework is designed to reconcile scenario execution engines, such as Scenic [2] and ScenarioRunner [20], and training frameworks in the AD research community. We model the interaction between the environment and the agents as a Partially Observable Stochastic Game (POSG) where all agents take actions simultaneously at each timestep. Thus, a natural choice for the environment interface is the PettingZoo API [21], the multi-agent version of OpenAI Gym [22].
III-A Architecture
Figure 2: MATS-Gym Architecture: We build on ScenarioRunner the scenario execution infrastructure. Blue: our contributions, Green: existing infrastructure from ScenarioRunner.
Figure 3: Action abstractions include low-level control, waypoints for local planning, and macro-actions for behavior planning.
We build MATS-Gym around the core of ScenarioRunner, which allows us to leverage their existing infrastructure to specify scenarios, define event observers and execute scenarios, as depicted in Figure 2. Our framework serves as an execution handler and training interface that extends the capabilities of those scenarios to multi-agent environments. Furthermore, we provide a set of wrappers to enhance observation and action spaces in order to facilitate versatile training environments. Additionally, we also provide a task infrastructure that allows to easily define reward functions and termination conditions. One of the design goals of MATS-Gym is interoperability with current scenario engines for CARLA. Compatibility with ScenarioRunner and thus OpenSCENARIO is inherently built in. As a consequence, our framework can be also used as a training and evaluation framework for the CARLA AD Challenge [^2] https://leaderboard.carla.org/challenge/. Furthermore, we connect another complementary scenario specification engine for CARLA by implementing an adapter for Scenic, a probabilistic programming language for building complex scenarios in various domains. Scenic comes with a rich set of building blocks for specifying distributions over traffic scenarios that adhere to predefined constraints.
III-B Observations, Actions and Tasks
MATS-Gym offers comprehensive infrastructure for agent state retrieval, providing ground truth information such as position, velocity vectors, and key traffic events like violations and collisions. Its modular design allows for flexible specification of additional observation types. It offers a versatile infrastructure for configuring sensor suites for agents, akin to the CARLA AD Challenge, supporting various sensors such as cameras, LiDAR, and Radar. It extends the CARLA BirdEyeView [23] framework for bird’s-eye view observations, constructing occupancy gridmaps around agents encoding lane markings, traffic lights, signs, and other road users. Agents can also access road network information and vector-based representations of the road layout. We account for different type of tasks by providing action spaces at various levels of abstraction. On the lowest level, agents can issue throttle, breaking and steering commands at high frequencies. On a higher level, agents can instead provide a target waypoint which is tracked by a PID controller for smooth driving trajectories. We also provide a discrete action space for high-level driving commands, suitable for route planning, behavior prediction, or multi-agent interaction. Fig. 3 visualizes the available action spaces. Furthermore, MATS-Gym allows for defining custom tasks and offers pre-defined tasks related to autonomous driving, including route following, infraction avoidance, and driving comfort, crucial for designing reward functions in reinforcement learning approaches.
IV Scenario Based Curriculum Generation
In the following, we will outline our approach to use Scenic’s parametrizable scenario specifications as a way to build an adaptive scenario generation procedure. We adopt a dual-curriculum design algorithm based on REPAIRED [3], a recent algorithm combining PLR with an adaptive level generator. At each timestep, we either generate a new level by sampling environment parameters from $\tilde{\pi}$ or we sample a replay level from our PLR buffer. The buffer samples levels with a probability proportional to their regret and a staleness metric, as described in the original approach. Then, we sample a trajectory from the environment to update the student policy $\pi$ and to compute the estimated regret for the current level, which is then used to update the buffer and the environment generator. Algorithm 1 outlines this approach.
Differently than the original, we do not use an antagonist agent to estimate the regret, due to computational constraints. Instead, we use the Maximum Monte Carlo [3] regret estimator, formally described as:
$\frac{1}{T}\sum_{t=0}^{T}R_{\text{max}}-V(s_{t}).$ (1)
This regret formulation compares the maximum achieved return with the average state value for each episode. We choose this alternative to avoid training another policy just for regret estimation, which would entail twice as much simulation steps. Furthermore, for the level generator optimization, we use the tree-structured Parzen estimator algorithm, implemented in the hyperparameter optimization framework Optuna [24] instead of RL.
Algorithm 1 MATS-REPAIRED
1:  Input: policy $\pi$, environment generator $\tilde{\pi}$, buffer $\Lambda$
2:  while not converged do
3:     Sample replay decision $d\sim P_{D}(d)$
4:     if d=0 then
5:        Generate environment parameters $\theta\sim\tilde{\pi}$.
6:        Insert $\theta$ into buffer $\Lambda$.
7:     else
8:        Sample replay level, $\theta\sim\Lambda$.
9:     end if
10:     Collect trajectory $\tau$ using policy $\pi$.
11:     Compute regret $R$ using Eq. 1.
12:     Update policy $\pi$ with $\tau$.
13:     Update level buffer: $\Lambda\leftarrow(\theta,R)$.
14:     if level was generated then
15:        Update environment generator $\tilde{\pi}$ with $R$.
16:     end if
17:  end while
V Experiments
We conduct two experiments in which we show that MATS-Gym can be used to (1) train multiple agents in an implicit coordination task and (2) for auto-curriculum generation. Both experiments share the base scenario in which agents are tasked with successfully navigating along a route through an intersection in urban environments, similar to the one introduced in the illustrative example. The reward function for this task is a combination of a progress-based and a cruise speed reward:
$r_{t}=r_{\text{progress}}+r_{\text{cruise}},$ (2)
where $r_{\text{progress}}=p_{t}-p_{t-1}$ is the progress made along the route since the last time-step and $r_{\text{cruise}}=min(\frac{v_{t}}{v_{\text{target}}},1)$ rewards the agent to drive at a cruise velocity $v_{target}$.
For the optimization of the agent policy we use Proximal Policy Optimization (PPO) [25]. The agent receives birdview observations as input and can therefore observe the drivable road area, lane markings, route and other traffic participants.
V-A Learning in Different Action Spaces
Figure 4: Learning curves for I-PPO under different action definitions and the impact on episodic return, collisions and route completion. Performance reports mean and standard deviation over 5 consecutive policy updates of the same run.
In this experiment, we demonstrate our framework’s usability for multi-agent training and discuss how the choice of actions affects the training of cooperative agents.
Scenario Description. Based on the initial scenario, we let four vehicles navigate through a four-way intersection, each with a designated route and a time limit to reach their destination. The starting conditions and goals vary per episode as outlined in Scenic. To succeed, the agent must manage its vehicle, collaborate with other traffic participants, and follow the prescribed route. Termination occurs when all agents are stopped.
Training Setup. Each vehicle is governed by an independent policy, which perceives a birdview observation encoding the road layout, vehicle positions relative to the agent, and route information. To study the impact of actions spaces, we consider the same task definition with continuous, waypoint and macro actions. Each of the proposed actions have different control frequencies: continuous actions are repeated twice, waypoint actions are executed for 5 steps, and macro actions persist for 10 steps with 0.05s per step. We train the policies with Independent PPO [26] and account for the different action frequencies by fixing the training budget to $175$ policy updates, each over a batch of $2048$ transitions.
Results. Figure 4 shows the learning curves with respect to average episodic reward, collisions and route completions, evaluated over the batch of training data for each policy update. We aggregate the metrics over $k$-size bins of policy updates ($k=5$), reporting the mean and standard deviation.
We observe different characteristics of the learned policies. Low-level and waypoint target actions incur a high number of collisions, especially in early training stages. On the other hand, macro actions make use of a basic lane keeping and collision avoidance controller, which prevents agents from leaving drivable areas and avoids most of the accidents, leading to a low collision rate throughout the training. This restriction comes at the cost of frequent deadlocks, which result in a low route completion rate. Less restrictive action spaces allow agents to leave the predefined lane to conduct evasive maneuvers, leading to higher average route completion in later training stages.
This experiment emphasizes how the action space design profoundly shapes emergent behavior and affects the learning task’s difficulty. Careful modelling of the problem, encompassing observations and actions, is crucial for multi-agent learning and serves as a key feature in training frameworks. Importantly, it demonstrates that there isn’t a universally optimal level of abstraction for the action space; instead, the choice depends on the problem at hand.
V-B Scenario Based Environment Design
Figure 5: Learning curves of average episodic return, route completion during training and evaluation over 3 seeds with one standard deviation. We also report the average regret of the level buffers of PLR and our DCD approach over timesteps.
In this experiment, we assess our DCD approach’s ability to align the training scenario distribution with the agent’s capabilities. We examine its impact on agent performance and the distribution of generated levels. Additionally, we compare its effectiveness with PLR and basic domain randomization. We analyze the evolution of sampled scenario parameters for each approach and evaluate their adaptation to the agent’s capabilities, particularly in the early stages of training when the agent struggles with complex scenarios.
Scenario Description. We parameterize the base scenario introduced before with a variety of discrete and continuous variables, which characterize the difficulty of the generated scenes. Intuitively, learning to follow a straight path through an empty intersection is easier than performing an unprotected left turn in a busy intersection where other traffic participants are driving recklessly. In our experiment, we define the following parameters that can be sampled:
  • Assigned route.
  • Number of other vehicles in the intersection.
  • Target speed for the other vehicles’ controllers.
  • Whether the other vehicles ignores others recklessly.
  • Whether the other vehicles respect traffic lights.
Results.
Figure 6: Evolution of the parameters that are sampled during the training with DCD (top) and domain randomization (bottom). For each parameter, we report mean and standard deviation over the batch of training data, indicate the direction of major difficulty ($\uparrow$), and the linear interpolation to highlight the general trend (red). We observe that DCD steers the parameter distribution towards configurations that are supposedly easier to solve: straight crossings, fewer and safer NPCs, etc. This leads us to the conclusion that DCD adapts the training distribution faster to the performance level of the agent.
In Figure 5, we compare learning curves over 150K environment steps. Both PLR and DR show similar performance in terms of episodic returns and route completion during training, indicating comparable scenario difficulty. DCD demonstrates a notable increase in training signal, suggesting adaptation by generating progressively easier scenarios in the initial stages of training. We evaluate all approaches on the same 12 hold-out levels, representing different maneuvers with varying numbers of NPCs. Although the performance on the hold-out set is comparable in early stages of the training, DCD and PLR do not suffer as much from performance drops in later stages.
Figure 6 provides insight into the progression of scenario parameters throughout training. DCD shifts the scenario distribution towards less challenging levels, characterized by straight maneuvers for the agent and a low number of vehicles in intersections. For comparison, we also show how DR naturally maintains a uniform distribution over environment parameters.
We also investigate the distribution of level parameters in the replay buffers of PLR and REPAIRED. Figure 7 depicts the average regret associated with various parameter combinations at four distinct checkpoints during the training process. For visualization purposes, we focus on combinations involving the maneuver type and the number of other vehicles (NPCs) present in the intersection. Observing the data, it becomes evident that the average regrets of Dual-Curriculum Design (DCD) levels tend to be higher and exhibit a narrower distribution across fewer parameter combinations compared to those of PLR alone. This trend can be attributed to the presence of the adaptive level sampler, which facilitates expedited convergence towards parameter combinations that are more pertinent to the task at hand.
Our experiments showcase the effectiveness of automatic curriculum design in aligning scenario generation with the agent’s capabilities. Additionally, optimizing the scenario generation process, rather than just the resampling procedure, accelerates the adaptation of the level distribution. Through this experiment, we aim to demonstrate that integrating UED approaches into MATS-Gym facilitates future research in curriculum generation for AD applications.
Figure 7: For each pair of maneuver type and number of NPCs, we average the regrets of all levels with the same parameters in the buffer at four checkpoints throughout the training. We observe that the environment generation policy of DCD leads to more narrowly concentrated regret on fewer level configurations, suggesting faster adaption of the level sampling.
VI Conclusion
In this work, we present MATS-Gym, a multi-agent training framework capable of generating scenario-based auto-curricula for AD tasks in CARLA. MATS-Gym is open to multiple, commonly used frameworks for traffic scenario specification and sampling, including ScenarioRunner, OpenSCENARIO and Scenic. Leveraging Scenic enables sampling from scenario distributions and the integration with UED approaches. This compatibility resembles a promising way to generate more relevant and realistic training scenarios for AD. We demonstrate the usability of our framework in two experiments of multi-agent training and automatic curriculum generation, respectively. By introducing MATS-Gym and demonstrating its application in various experiments, we contribute to ongoing efforts in advancing multi-agent training for autonomous driving.
For future work, we intend to explore the potential of MATS-Gym in adversarial training within multi-agent systems and the integration of generative models for traffic scenario generation.